# DiVerify: Hardening Identity-Based Software Signing with Programmable Diverse-Context Scopes

### v2

@@ Abstract
Code signing enables software developers to digitally sign their code using cryptographic keys, thereby associating the code with a specific key. This key is then linked to an identity (e.g., through an identity provider), allowing users to establish trust in the origin of the signature and verify both the code’s origin and integrity. However, this code–identity binding is only as trustworthy as the mechanisms enforcing it. State-of-the-art identity-based code signing schemes have a major shortcoming: they fail to provide verifiable information about the context in which a signature is generated. If an identity verification server is compromised or the signing client behaves maliciously, the resulting signature may falsely suggest a trustworthy origin, despite the absence of actual developer intent.
To address these issues, we propose a diverse identity verification approach that reduces reliance on a single source of verification and enforces stronger guarantees around the signing process itself. By combining multiple identity signals with verifiable execution environments, our system improves confidence that signatures reflect the intent of a legitimate user, produced under expected conditions. Signing in our DiVerify prototype incurs only a few kilobytes of additional storage (less than 0.4% of the average package size in widely used ecosystems like PyPI) In DiVerify, signing completes in under 100ms on a typical deployment.
“Yo soy yo y mi circunstancia; y si no la salvo a ella no me salvo yo.”111“I am I and my circumstance; and, if I do not save it, I do not save myself.” — José Ortega y Gasset
1 Introduction
@@
Software supply chain attacks have become a concern due to their large-scale impact on software users’ safety [73, 71, 9] These attacks involve the injection of malicious code into a software artifact and subsequent exploitation in downstream systems [64]. In response, government and industry groups have begun to advocate for improved provenance along the software supply chain [13, 26, 20].
@@
The most common provenance method for software supply chains is software signing. Software providers attach a digital signature to an artifact using public key cryptography, and software consumers can then test this signature to detect tampering [14]. Despite decades of use [74, 19], adoption has been limited by usability [35] and key management challenges [79, 36]. To address this, identity-based signing platforms emerged [78], still grounded in public key cryptography [62, 49], but linking signatures to verified identities via identity providers (IdPs) and short-lived ephemeral keys to allow a single signing event.
@@
However, the state-of-the-art identity-based signing solutions are not resilient. If any party involved in the signing operation is compromised, then the provenance guarantees are lost. This limitation manifests in two ways. First, identity-based signing systems rely on a single IdP to identify a signer and issue authentication tokens used for signing. If this identity provider is compromised, it renders all associated signing events untrustworthy. Many recent attacks have targeted identity providers, including the Midnight Blizzard intrusion [55], Apple’s SSO [6], and attacks on Facebook [69, 76].
The second limitation is the signing client, which brokers interactions between the software author, the identity provider, and the repository. These clients are entrusted with signing capabilities for arbitrary code during the validity of their ephemeral key. In this case, the security of code signing on the client side relies on the effectiveness of client-side security measures. If compromised, they can sign malicious software that appears fully legitimate. An example is the attack by the North Korea-based threat actor Diamond Sleet (ZINC) [54], where attackers exploited a code-signing system’s privileges to sign malicious variants of an application developed by CyberLink, making it appear legitimate. Current solutions rely on retrospective monitoring of signature logs to detect unauthorized signing [62], which only occurs once malicious software has been distributed and utilized.
These two problems share a common cause: inadequate transparency and verifiability of the conditions under which a signature was produced. Current systems provide no mechanisms for verifiers to assess the signer’s security posture or signing context, creating a blind spot that attackers can exploit using valid credentials under malicious conditions. To address this, software signing systems must evolve beyond verifying who signed a payload, and must also verifiably communicate how a signature was produced – under what conditions, on what systems, and with what constraints. In other words, we need mechanisms to make the signing context itself verifiable, so that trust is no longer concentrated in a static identity or a privileged client but distributed across a transparent, auditable chain of evidence.
Towards this goal, in this paper we present DiVerify, a system that enhances identity-based signing by enforcing a verifiable, contextual security posture at signing time. DiVerify introduces a programmable scope verification framework that expands the notion of signer identity to include a flexible set of verifiable attributes – called DiVerify scopes. DiVerify operates under a zero-trust assumption toward signing clients. Rather than assuming the integrity of the client environment, it requires that all sensitive signing operations occur within a hardware-protected Trusted Execution Environment (TEE). Each signing operation produces a DiVerify proof: a cryptographically verifiable attestation of both identity attributes and the runtime integrity of the signing environment. This enables the enforcement of fine-grained, context-aware policies for signature verification, reducing implicit trust and raising the bar for a successful client compromise.
The core novelty of DiVerify is that it extends identity-based code signing to enforce how software is signed, not just by whom. DiVerify enables the enforcement of multi-independent identity and client attestations at signing time, ensuring verifiable secure context for software signatures.
@@
We evaluate DiVerify in terms of its security guarantees and operational overhead. To assess security, we show how different DiVerify instantiations can balance security and performance across different variants to ease integration into existing identity-based signing systems, and highlight how they mitigate the adversaries described in our threat model. To assess overhead, we quantify the performance and storage costs of signing and verification operations across the different variants. To demonstrate its practical relevance, we analyze several historical software supply chain attacks and illustrate how DiVerify’s model could have helped detect or prevent them. Our evaluation shows that DiVerify introduces minimal overhead compared to current identity-based software signing solutions with signing operations complete in under 100ms for regular deployment, and verification adds only  11% latency compared to legacy-compatible approaches, highlighting its feasibility for real-world adoption.
@@
  1. 1.
    A novel framework for context-aware identity verification in code signing, which, for the first time, combines multiple identity proofs with hardware attestation.
  2. 2.
    Design and implementation of DiVerify, a prototype that enforces context-bound signing policies using TEE-based attestations, an approach not present in prior signing tools.
  3. 3.
    Evaluation of security and performance under different settings, demonstrating that enforcing multi-identity verification with TEE-backed signing is practical, which was previously unexamined.
  4. 4.
@@
    Security analysis showing that DiVerify would have prevented several real-world supply-chain attacks that prevailing solutions could not have.
2 Background
@@
To describe DiVerify, we first discuss its context (i.e., the software supply chain) and its role (i.e., code signing). The DiVerify protocol also requires understanding authentication mechanisms used to establish identity and context, and remote attestation, which we leverage in our implementation.
2.1 Software Supply Chain and Code Signing
@@
The software supply chain involves all the activities carried out in the task of creating and distributing software, such as coding, compiling, testing, packaging and deployment [79, 64]. These activities rely on a wide range of tools, libraries, frameworks, and services creating a complex web of dependencies [42, 83][42, 83]. This complexity makes it a prime target for attackers, as seen in incidents like SolarWinds [24] and Log4j [60]. As a result, organizations are increasingly adopting practices like code signing, improved dependency management and vetting [68, 18], continuous monitoring [84], zero-trust architectures [46, 86, 1], and Software Bills of Materials (SBOM) [61] to ensure transparency, integrity, and compartmentalization throughout the supply chain.
@@
Code signing allows a verifier to establish trust with a software producer. The process uses a keypair $(k_{p},k_{s})$ and an artifact $(SW_{a})$ to generate a signature $(S_{sw})$. A verifier will then use $(K_{p},SW_{a},S_{sw})$ to verify whether $(SW_{a})$ is, first, integral (i.e., hasn’t been tampered with), and second, belongs to a trusted party. Traditional tools like OpenPGP [67] use manual key management and user-controlled trust models [8, 87], requiring developers to generate, safeguard, and distribute long-lived keys, introducing usability challenges and risks of compromise. Identity-based solutions such as Sigstore [62], OpenPubKey [25], and Keyfactor [38] simplify this by using ephemeral keys and identity federation protocols (e.g., OIDC, OAuth, SAML). These allow developers to prove identity ownership via standard authentication flows, then linking identities to signing keys through certificates or embedded tokens.
2.2 Authentication Mechanisms
@@
Identity-based signing tool uses authentication to verify the identity of a signer. Authentication mechanisms vary, but generally fall into categories based on the type of claim they make [27, 51]: identity, possession, or context.
@@
Identity claims verify who a signer is, using protocols like OAuth 2.0 [63] and OpenID Connect [65]. These protocols issue signed tokens (e.g., JWTs [4]) from trusted providers like Google. Distributed Identities (DiD) [89] and CryptID [28] provide mechanisms similar to OIDC but act as a meta-protocol enabling identity-specific claims.
@@
Possession claims prove control over physical elements such as a hardware token, smartcard, or cryptographic key. The authenticated party demonstrates possession of a secret (e.g., a private key) through a cryptographic challenge-response protocol. Protocols like FIDO2/WebAuthn, use devices (e.g., YubiKeys [91]) to protect the private key and perform challenge responses locally ensuring that only someone with the device can authenticate.
@@
Contextual claims assert properties of the environment in which events occur, focusing on system state rather than identity or key possession. They support trust decisions based on runtime conditions like device or software integrity. Techniques include device fingerprinting [75], TPM-based measurements [53, 32], and kernel-level attestation via Linux’s Integrity Measurement Architecture (IMA) [48]. In DiVerify, such claims help bind events to execution contexts.
2.3 Trusted Execution Environments
@@
Trusted Execution Environments (TEEs) isolate processes using hardware mechanisms like secure memory and privileged execution contexts [3]. Remote Attestation enables verifiers (e.g., a remote user or system) to confirm software execution within a secure and trusted environment on a remote machine. This is achieved through a cryptographic protocol in which the remote system generates a signed statement known as an attestation report or quote that includes a measurement (e.g., hash) of the software and configuration loaded into a secure execution environment. The verifier can compare this measurement against a known trusted value to determine whether the software executing is indeed the expected one.
@@
Trust in remote attestation relies on hardware-rooted chains of trust, such as those provided by Intel Software Guard Extensions (SGX) [17], Trusted Domain Extensions (TDX) [31] which include embedded attestation keys and associated endorsement certificates issued by the hardware manufacturer. These cryptographic materials ensure that only legitimate, manufacturer-approved hardware can produce valid attestation signatures.
3 Parties, Roles, and Threat Model
@@
Next we formalize identity-based software signing systems, the context in which they operate, and the security goals of both identity-based software signing in general, and DiVerify in particular.
3.1 Parties and Roles
@@
We consider a standard identity-based software signing setting with the following parties interacting with each other:
@@
Signer: an entity (i.e., a developer or an automated system) who authored the software artifact and intends to generate a signature $(S_{s}w)$ for it.
@@
Identity Provider (IdP): an entity that authenticates the Signer through the Client and provides qualified attributes about the User. In DiVerify we rely on a general-purpose instantiation of an IdP we call scope providers as described in section˜5.3.
@@
Package repository: a storage location where software packages and their signatures are published, shared, and retrieved for distribution and verification, e.g., NPM or Hugging Face.
@@
Verifier: The counterpart of the Signer. It verifies the signed software to establish trust by validating a signature against a software artifact using a certificate or public key.
@@
Figure 1: State-of-art identity-based code signing system model, highlighting the types of attacker capabilities that such systems are not resilient against. These attack types are discussed in section˜3.2.
@@
A typical high-level interaction flow is illustrated in Figure˜1. There are five steps. 1. The signer initiates a signing request by interacting with a client. 2. The client redirects her to an IdP to authenticate. 3. The signer authenticates, an identity token is obtained, and sent to the client. 4. Optionally, the client requests a signing certificate (which binds the identity to a signing key) from a certificate authority (or CA). 5. Finally, the client signs the software and uploads the signature to a repository.
A successful flow allows a verifier to obtain a $(S_{sw},K_{p},SW_{a})$ tuple from a source (e.g., a repository) and ensure the artifact is not tampered with.
3.2 Threat Model
Attackers intend to subvert the flow described above to deliver tampered software to verifiers (i.e., software users). In order to do so, they will impersonate, tamper with, and/or attack all the roles described above. A successful attack is one that causes a verifier to accept a malicious artifact $(SW_{a})$.
At a fundamental level, we want to ensure code signatures accurately link to the legitimate Signer’s identity and reflect their intent. Hence, we consider threats where adversaries seek to produce valid signatures without the consent of the legitimate identity owner.

### v3

# DiVerify: Hardening Identity-Based Software Signing with Diverse-Context Scopes

@@ Abstract.
Identity-based code signing enables software developers to digitally sign their code using cryptographic keys. This key is then linked to an identity (e.g., through an identity provider), allowing signers to verify both the code’s origin and integrity. However, this code–identity binding is only as trustworthy as the mechanisms enforcing it. State-of-the-art identity-based code signing schemes present a major shortcoming: these schemes fail to provide verifiable information about the context in which a signature is generated. This verifiability is crucial given that modern attackers have subverted long-established security assumptions, namely, that the identity provider ecosystem, as well as signing software itself is trusted.
To address these issues, this paper introduces a diverse identity verification framework, DiVerify, that distributes identity-based verification across multiple entities and enforces stronger guarantees about the signing context. DiVerify makes it possible to provide end-to-end verifiability of not only a signer’s identity (via multiple such signals), but also a signer’s software stack (e.g., to verify no malware is present on a system at the time of signing). DiVerify is aimed at deployability, and leverages a meta-protocol to gather various trust signals and a binding mechanism to address the aforementioned, novel software supply chain attack vectors. We evaluate DiVerify’s performance and confirm it is cheap to deploy and non-intrusive to developers: it only incurs a few kilobytes of additional storage (less than 0.4% of the average package size in widely used ecosystems like PyPI), and signing completes in under 100ms on a server-grade deployment.
“Yo soy yo y mi circunstancia.”
“I am I and my circumstance.”
— José Ortega y Gasset
Software Signing, Software Supply Chain Security, Policy Transparency
$c$s: Security and privacy Trust frameworksccs: Security and privacy Software security engineeringccs: Security and privacy Distributed systems security
1. Introduction
@@
Software supply chain attacks are a major, widespread concern due to their large-scale impact on software users’ safety (Reuters, 2021; Rapid7, 2021; CNCF TAG-Security, 0 12). These attacks involve the injection of malicious code into a software artifact and subsequent exploitation in downstream systems (Okafor et al., 2022). In response, government, industry and academia have proposed means to improve provenance along the software supply chain (Cooper et al., 2018a; House, 2021; Foundation, 2021).
@@
The most common provenance mechanism for in this setting is software signing (Rivest et al., 1978; Kalu and Davis, 2025). Software providers attach a digital signature to an artifact using public key cryptography, and software consumers can then test this signature to detect tampering (Cooper et al., 2018b). Despite decades of use (Rivest et al., 1978; Diffie and Hellman, 2022), adoption has been limited by usability (Whitten and Tygar, 1999; Kalu et al., 2025a) and key management challenges (Schorlemmer et al., 2024; Kalu et al., 2025b). To address this, identity-based signing platforms have emerged (Schorlemmer et al., 2025), still grounded in public key cryptography (Newman et al., 2022; Lorenc, 2023), but linking signatures to verified identities via identity providers (IdPs) and short-lived ephemeral keys to allow a single signing event.
@@
However, the state-of-the-art identity-based signing solutions are not resilient to compromises within their underlying trust model. If any party involved in the signing operation is compromised, then the provenance guarantees are lost. This limitation manifests in two ways. First, identity-based signing systems rely on a single IdP to identify a signer and issue authentication tokens used for signing – if this IdP is compromised, it renders all associated signing events untrustworthy. Many recent attacks have targeted identity providers, including the Midnight Blizzard intrusion (Microsoft, 2024), Apple’s SSO (Jain, 2020), and attacks on Facebook (Poza, 2018; Rosen, 2018). Second, signing clients responsible for brokering interactions between the software author, the identity provider, and the repository hold broad authority during the validity period of their ephemeral keys. If compromised, they can be coerced into producing valid signatures over malicious software. The Diamond Sleet (ZINC) attack (Microsoft, 2023), which abused code-signing system’s privileges to sign malicious variants of an application, exemplifies this risk. Current systems make an undue assumption that both Identity providers and clients will never be compromised.
In order to surmount this limitation, identity-based signing must provide verifiable evidence regarding the context in which signing took place. Verifiers should confirm how the signer authenticated or whether the signing client is not compromised. This capability addresses both problems outlined above. In sum, by forwarding verifiable evidence about the security posture of the signing process, verifiers can ensure that the signer is trusted, and that the signing event was carried out in a way that can be trusted as well.
Although many ad-hoc mechanisms exist to harden identity verification such as multi-factor authentication (Golla et al., 2021; Reese et al., 2019), hardware-backed credentials (Yubico, [n.d.]; Alliance, 2022), device-binding schemes (Hodges et al., 2021), and even cross-device secret-sharing approaches (Laing et al., 2022), state-of-the-art identity-based signing systems do not provide verifiers with verifiable evidence of these conditions. Similarly, client integrity can be reinforced through attestation (Kim et al., 2015) or other local hardening techniques (Wilkins and Richardson, 2013), these assurances never reach the verifier. And without evidence, the verifier cannot distinguish a legitimate signing activity from one carried out on a compromised client, leaving attackers free to exploit valid credentials under malicious conditions. We posit that this ability to transparently transfer this information to the verifier is crucial to addressing supply chain attacks targeting trust infrastructure (Laurie, 2014).
In this paper we present DiVerify, a framework for identity-based signing that addresses these challenges, and a prototype system that instantiates it. DiVerify extends existing system threat model and introduces two main components. The first is the DiVerify proof, a cryptographically verifiable bundle produced for each signing event that binds together identity attributes, authentication context, and attested measurements of the signing environment. Our construction provides a meta-protocol that integrates common, reliable mechanisms. Our meta-protocol combines these heterogeneous signals into a homogeneous proof that exposes independently-verifiable trust assertions. A second component, a DiVerify policy, specifies the conditions under which a verifier should accept a proof — e.g., dictating the minimum authentication strength, acceptable identity claims, and expected trusted execution environment integrity — and enables these conditions to be checked independently of the upstream signing pipeline. These policy constructions are essential to accommodate developer signing practices and provide expressions that can integrate verifiability for these heterogeneous signals. Together, these allow verifiers to enforce explicit, context-aware trust requirements rather than inheriting implicit assumptions identity providers or signing clients make.
The design of Diverify accounts for providing verifiability even as trust expectations change. Developers replace devices, add authentication factors, or migrate across platforms, and maintainers may strengthen or revoke scopes, making communication shift in DiVerify policies is paramount. Thus, DiVerify introduces a principled mechanism for representing and evolving verifier trust in a controlled and verifiable manner, ensuring that trust decisions remain consistent even in an adversarial environment.
@@
We evaluate DiVerify against the adversary in our threat model and show that DiVerify strengthens the security posture of the systems by increasing the adversarial effort required for compromise. We also examine how different DiVerify instantiations balance security and performance trade-offs when integrated into existing signing systems. To demonstrate practical relevance, we analyze several historical supply-chain attacks and show that DiVerify could have detected or prevented them. Finally, we measure the performance overhead of our prototype and find that signing completes in under 100 ms and verification adds only 11% latency compared to legacy-compatible approaches, highlighting its feasibility for real-world adoption.
@@
  1. (1)
    Based on recent attacks, we extend the existing identity-based signing system threat model. We then propose a novel verifier-centric trust model for identity-based code signing. We treat signing context, identity, authentication, and signing-environment integrity as a single, verifiable proof.
  2. (2)
    We develop mechanisms for authenticated, evolvable policy state that allow verifiers to detect replayed or downgraded trust configurations and reason about policy-level attacks not covered by existing identity-based signing systems.
  3. (3)
    We design and implement DiVerify, a system that enforces context-aware signing policies over these proofs, replacing ad hoc verifier configuration with an explicit policy model.
  4. (4)
@@
    We evaluate DiVerify’s security and performance, measuring the cost of generating and verifying proofs and showing that DiVerify can mitigate several real-world attacks that prevailing solutions could not.
Significance
Identity-based software signing has become a central security primitive across modern software ecosystems, providing the attribution signal upon which downstream trust decisions depend. This model underlies workflows in cloud platforms (The Kubernetes Authors, 2 11), AI and MLOps model pipelines (Sablotny, 2025; Font, 2025), and package ecosystems (npm, Inc., 5 22; Ingram, 2024). DiVerify reinforces this primitive by reducing reliance on any single authority and by mitigating compromise of the signing primitive itself. Although we instantiate DiVerify for software signing, the underlying model generalizes to any system where provable attribution is the foundational security property upon which all downstream trust relationships depend.
2. Background
@@
To describe DiVerify, we first discuss its intended role: code signing in the software supply chain (section 2.1). The DiVerify protocol also requires understanding authentication mechanisms (section 2.2) used to establish identity and context, and remote attestation, which we leverage in our implementation.
2.1. Software Supply Chain and Code Signing
@@
The software supply chain involves all the activities carried out in the task of creating and distributing software, such as coding, compiling, testing, packaging and deployment (Schorlemmer et al., 2024; Okafor et al., 2022). These activities rely on a wide range of tools, libraries, frameworks, and services, creating a complex web of dependencies (Kumar et al., 2017; Synopsys, 2023a; Kumar et al., 2017; Synopsys, 2023a). This complexity makes it a prime target for attackers, as seen in incidents like SolarWinds (Google Cloud, 2020) and Log4j (National Cyber Security Centre, 2021). As a result, organizations are increasingly adopting practices like code signing, improved dependency management and vetting (Williams et al., 2025; Pashchenko et al., 2020; de Souza and Redmiles, 2008), continuous monitoring (Synopsys, 2023b), zero-trust architectures (Lamb and Zacchiroli, 2022; Torres-Arias et al., 2019; Amusuo et al., 2025), and Software Bills of Materials (SBOM) (National Telecommunications and Information Administration, 2021; Abu Ishgair et al., 2025) to ensure transparency, integrity, and compartmentalization throughout the supply chain.
@@
Code signing allows a verifier to establish trust with a software producer (Rivest et al., 1978; Kalu and Davis, 2025). The process uses a keypair $(k_{p},k_{s})$ and an artifact $(SW_{a})$ to generate a signature $(S_{sw})$. A verifier will then use $(K_{p},SW_{a},S_{sw})$ to verify whether $(SW_{a})$ is, first, integral (i.e., hasn’t been tampered with), and second, belongs to a trusted party. Traditional tools like OpenPGP (OpenPGP Project, [n.d.]) use manual key management and user-controlled trust models (Chandramouli et al., 2013; Unger et al., 2015), requiring developers to generate, safeguard, and distribute long-lived keys, introducing usability challenges and risks of compromise. Identity-based solutions such as Sigstore (Newman et al., 2022), OpenPubKey (Heilman et al., 2023), and Keyfactor (Keyfactor Team, 2025) simplify this by using ephemeral keys and identity federation protocols (e.g., OIDC, OAuth, SAML). These allow developers to prove identity ownership via standard authentication flows, then linking identities to signing keys through certificates or embedded tokens.
2.2. Authentication Mechanisms
@@
Identity-based signing uses authentication to verify the identity of a signer. Authentication mechanisms vary, but generally fall into three categories (Huang et al., 2010; Mahadi et al., 2018): identity, possession, or context.
@@
Identity claims verify who a signer is, using protocols like OAuth 2.0 (OAuth.net, 2 05) and OpenID Connect (OpenID Foundation, 2023). These protocols issue signed tokens (e.g., JWTs (Jones et al., 2015)) from trusted providers like Google. Distributed Identities (DiD) (World Wide Web Consortium, 2022) and CryptID (Cryptid, [n.d.]) provide mechanisms similar to OIDC but act as a meta-protocol for identity-specific claims.
@@
Possession claims prove control over physical elements such as a hardware token, smartcard, or cryptographic key. The authenticated party demonstrates possession of a secret (e.g., a private key) through a cryptographic challenge-response protocol. Protocols like FIDO2/WebAuthn, use devices (e.g., YubiKeys (Yubico, [n.d.])) to protect the private key and perform challenge responses locally ensuring that only someone with the device can authenticate.
@@
Contextual claims assert properties of the execution environment, focusing on system state rather than identity or key possession. They support trust decisions using runtime conditions like device or software integrity. Techniques include device fingerprinting (Rizzo et al., 2021), TPM-based measurements (Microsoft, 4 21; Intel Corporation, 2025), and kernel-level attestation via Linux’s Integrity Measurement Architecture (IMA) (Linux Integrity Subsystem, 4 21).
DiVerify extends identity-based signing by supporting possession and contextual claims in addition to identity claims, allowing verifiers to incorporate evidence about the signing environment and execution context into acceptance decisions.
2.3. Trusted Execution Environments
@@
Remote Attestation enables verifiers (e.g., a remote user or system) to confirm software execution within a secure and trusted environment on a remote machine. Trusted Execution Environments (TEEs) enable this capability by isolating processes using hardware mechanisms like secure memory and privileged execution contexts (Arm Ltd., 4 21).
This is achieved through a cryptographic protocol in which the remote system generates a signed attestation report or quote that includes a measurement (e.g., hash) of the software and configuration loaded into a secure execution environment. The verifier can compare this measurement against a known good value and determine if the software executing is indeed the expected one.
@@
Trust in remote attestation relies on hardware-rooted chains of trust, such as those provided by Intel Software Guard Extensions (SGX) (Costan and Devadas, 2016), or Trusted Domain Extensions (TDX) (Intel Corporation, 2021) which include embedded attestation keys and associated endorsement certificates issued by the hardware manufacturer. These cryptographic materials ensure that only legitimate, manufacturer-approved hardware can produce valid attestation signatures.
3. Parties, Roles, and Threat Model
@@
We formalize identity-based software signing systems and the context in which they operate (section 3.1). We show that existing threat models are too weak relative to recent attacker behavior and define the security goals of both identity-based software signing in general, and DiVerify in particular (section 3.2).
3.1. Parties and Roles
@@
We consider a standard (Grassi et al., 2017) identity-based software signing setting with the following parties interacting with each other:
@@
Signer: an entity (i.e., a developer or an automated system) who intends to generate a signature $(S_{s}w)$ over an artifact.
@@
Identity Provider (IdP): an entity that authenticates the Signer through the Client and provides qualified attributes about the User. In DiVerify we rely on a general-purpose instantiation of an IdP we call scope providers as described in section 5.3.
@@
Package registry: a storage service for software packages and their signatures. Users publish, share, and retrieve artifacts for distribution and verification, e.g., NPM (JS) or Hugging Face (ML).
@@
Verifier: is the counterpart of the Signer. It ensures that the signed software is valid by validating a signature against a software artifact using a certificate or public key.
@@
Figure 1. State-of-the-art identity-based code signing system model, highlighting the types of attacker capabilities that such systems are not resilient against. These attack types are discussed in section 3.2.
@@
A typical high-level interaction flow is illustrated in Figure 1. There are five steps. (1) The signer initiates a signing request by interacting with a client. (2) The client redirects her to an IdP to authenticate. (3) The signer authenticates, an identity token is obtained, and sent to the client. (4) Optionally, the client requests a signing certificate (which binds the identity to a signing key) from a certificate authority (or CA). (5) Finally, the client signs the software and uploads the signature to a registry.
A successful flow allows a verifier to obtain a $(S_{sw},K_{p},SW_{a})$ tuple from a source (e.g., a registry) and ensure the artifact is not tampered with. The verifier must also evaluate the identity attributes associated with the public key $K_{p}$ such as the signer’s email address and the issuing identity provider against locally configured trust expectations. These identity checks determine whether the signature corresponds to an authorized signer and thus whether the artifact should be accepted.
3.2. Threat Model
We consider a stronger threat model than previous literature to reflect modern attacker behavior. Components traditionally treated as trusted are now realistic targets whose compromise directly undermines security guarantees.
Prior work on the software supply chain has modeled threats to specific segments of the chain, such as source-code integrity (Vu et al., 2021; Gonzalez et al., 2021), build-system compromise (Fourné et al., 2023; Ohm et al., 2020), maintainer account compromise (Zimmermann et al., 2019; Samuel et al., 2010; Gu et al., 2023; Huang et al., 2024), package registry (Kuppusamy et al., 2016; Zimmermann et al., 2019; Zahan et al., 2022; Wu et al., 2023; Nikitin et al., 2017), mirror compromise (Cappos et al., 2008), package-distribution (Gao et al., 2024) and dependency-resolution attacks (Neupane et al., 2023; Vu et al., 2020). These models assume a trusted identity provider and signing client and rely on log transparency (Al-Bassam and Meiklejohn, 2018; Blauzvern, 2023; Laurie, 2014) to detect misbehavior after an event has been maliciously signed.
For instance, Log Transparency, places the responsibility for detecting misuse on signers, while the security consequences fall on downstream users. Because detection is retrospective, any delay between compromise and discovery exposes verifiers to the risk of accepting malicious artifacts. Yet verifiers must make trust decisions at verification time, and current systems provide them with no direct evidence about whether a signature reflects legitimate authentication or signer intent.
Recent incidents have shown that these components are realistic targets for attackers (Microsoft, 2024; Jain, 2020; Poza, 2018; Rosen, 2018). Excluding these capabilities yields a threat model that is too weak to capture the risks faced by modern software supply chains.
This motivates a threat model in which verifiers can validate these conditions themselves, rather than depend solely on retrospective mechanisms that upstream parties may not notice or act upon. We therefore extend prior threat models to treat the identity and signing ecosystem as an attack surface, capturing adversaries who can produce valid signatures without the legitimate signer’s intent. Regardless of their capabilities, all threat models in this domain assume that attackers aim to cause verifiers to accept tampered software. To do so, they may impersonate, tamper with, and/or attack any of the roles in the signing workflow, or manipulate the verifier’s locally configured trust expectations.
Attacker Goal
Ultimately, for all these threat models, attackers are successful if they can cause a verifier to accept a malicious artifact $(SW_{a})$. We want to raise the cost for such compromises by increasing the evidence an attacker must produce to induce acceptance.